A sample of 25 freshman nursing students made a mean score of 77 on a


test designed to measure


attitude toward the dying patient. The sample standard deviation was 10. Do


these data provide


sufficient evidence to indicate, at the .05 level of significance, that the


population mean is less than


80? What assumptions are necessary?

1
Expert's answer
2022-03-22T19:16:03-0400

"H_0:\mu \geqslant 80\\H_1:\mu <80\\T=\sqrt{n}\frac{\bar{x}-\mu}{s}=\sqrt{25}\frac{77-80}{10}=-1.5\sim t_{n-1}=t_{24}\\P-value:\\P\left( T<-1.5 \right) =F_{t,24}\left( -1.5 \right) =0.0733"

Since the P-value exceeds the significance level, there is not enough evidence that the mean value is less than 80.

The assumption is normality.